# IOI: Invisible One-Iteration Adversarial Attack on No-Reference Image- and Video-Quality Metrics

### v1

@@ Abstract
No-reference image- and video-quality metrics are widely used in video processing benchmarks. The robustness of learning-based metrics under video attacks has not been widely studied. In addition to having success, attacks on metrics that can be employed in video processing benchmarks must be fast and imperceptible. This paper introduces an Invisible One-Iteration (IOI) adversarial attack on no-reference image and video quality metrics. The proposed method uses two modules to ensure high visual quality and temporal stability of adversarial videos and runs for one iteration, which makes it fast. We compared our method alongside eight prior approaches using image and video datasets via objective and subjective tests. Our method exhibited superior visual quality across various attacked metric architectures while maintaining comparable attack success and speed. We made the code available on GitHub.
@@ 1 Introduction
NR metrics are employed in various image- and video-processing benchmarks, such as super-resolution (isr, 2019; Khrulkov & Babenko, 2021; 10., 2017), video generation (vid, 2024), video compression (Ghadiyaram et al., 2017). For some tasks, NR metrics show even better performance than FR metrics; for example, video super-resolution (sr-, 2023) or video compression by new encoding standards such as H.266/VVC (Antsiferova et al., 2022). The developers of video-processing algorithms can integrate adversarial attacks on quality metrics into their methods to achieve higher positions in public benchmarks. Today, such cheating can be detected in benchmarks that publish subjective comparisons along with objective ones. For example, in MSU Subjective Comparison 2021 (msu, 2021), the leaderboard by a learning-based metric VMAF, which was shown to be vulnerable to attacks (Siniukov et al., 2022; Zvezdakova et al., 2019), differs from a subjective one. Algorithms in these benchmarks compete in both visual quality and speed, and a high speed is essential for some real-life applications like universal encoding for video compression (from one to ten frames per second). Thus, to cheat in video processing benchmarks, it is profitable for an attacker to inject an imperceptible perturbation into the video without significantly decreasing the method speed.
@@
    High speed of an attack. The attack must operate at high speed for practical viability. A slow attack holds limited practical significance as its integration into video processing algorithms will greatly slow down the method and lower its position in benchmarks.
@@
Our research aims to investigate the potential of injecting fast, invisible and temporally consistent adversarial attacks on NR quality metrics. This paper introduces the Invisible One-Iteration (IOI) adversarial attack for images and videos. To achieve high attack speed, our method yields perturbation by calculating the gradient of an attacked model using one access to the model. We further show that a one-iteration attack at each frame is more efficient than a many-iteration attack applied to not every frame. To keep the temporal stability of a perturbed video, we make our attack invisible to the human eye by using weighting and frequency modules. The proposed attack operates in a white-box scenario, which does not limit its applicability for benchmarks: usually, a comparison methodology is known, and quality metrics are published for reproducibility. The primary contributions of this work can be summarized as follows:
@@
Our code is publicly available at link is hidden for a blind review.
@@ 2 Related Work
Numerous studies showed that $L_{p}$ norms are not suitable as a distance metric to evaluate perceptual image quality (Sharif et al., 2018; Fezza et al., 2019; Wang et al., 2004; Johnson et al., 2016; Isola et al., 2017). Perturbations in images generated under $L_{p}$ norm constraints often result in noisy pixels within smooth areas of the original image, which is easily perceptible to the human eye. Several adversarial attacks prioritize the visual quality of the generated adversarial images and use more sophisticated restrictions on perturbations than the bare implementation of $L_{p}$ constraints. AdvJND (Zhang et al., 2020) introduces a novel approach by incorporating just noticeable difference (JND) (Yang et al., 2005) coefficients into the $L_{\infty}$ norm constraint during adversarial example generation. These coefficients account for the human eye’s ability to perceive the threshold of changes in an image. The authors employed I-FGSM and FGSM algorithms as baselines. To enhance the visual fidelity of adversarial images, they amplified original perturbations obtained from the FGSM or I-FGSM methods by scaled JND coefficients. SSAH (Luo et al., 2022) adversarial attack targets two objectives: semantic similarity of images and low-frequency constraint. The first component is crafted with a focus on image classification tasks. The second component adds perturbations within high-frequency regions by minimizing the difference between low-frequency information of adversarial and original images. Usually, it requires many iterations to converge and produce good visual quality. Korhonen et al. (Korhonen & You, 2022b) introduced an iterative attack method targeting NR quality metrics, employing a Sobel filter to hide distortions within textured regions. In each iteration, the model’s gradient under attack concerning the input is multiplied by a spatial activity map derived from the original image via the Sobel filter. This map highlights areas with substantial texture, enhancing the visual fidelity of the resulting adversarial images. Zhang et al. (Zhang et al., 2022a) introduced an iterative adversarial image crafting approach leveraging various FR metrics like Chebyshev distance, SSIM, LPIPS, and DISTS to manage visual distortions. Their method involved the iterative minimization of a loss function consisting of two components: the attacked loss and a loss based on some differential FR image quality metric. Karli et al. introduced the Normalized Variance Weighting (NVW) (Karli et al., 2021) method aimed at amplifying perturbations within high-variance regions of images. This technique can be utilized alongside gradient attacks like FGSM or I-FGSM. Additionally, the authors proposed the LPIPS-minimization method, aiming to enhance perceptual quality by minimizing the LPIPS distance between the original and adversarial images while ensuring the classifier remains deceived. However, this minimization method is exclusively effective for discrete tasks, such as classification or detection, while applying it to quality metrics will lead to eliminating relative gain.
@@ 3.1 Problem formulation
The adversarial attack on the NR quality metric $M$ is modelled as follows:
@@
To ensure a high visual quality of perturbed image, instead of using $l_{p}$ norms that are inefficient for this task, we formulate the problem as follows:
$\underset{I^{a}}{\mathrm{argmax}}\{M(I^{a})-M(I)\},\frac{L_{f}(I^{a})\cdot L_{f}(I)}{||L_{f}(I^{a})||\text{ }||L_{f}(I)||}\leq\epsilon,$ (2)
where $L_{f}$ is a low-frequency filter. When low-frequency components of the clear and adversarial images are closely aligned, the perturbations mainly affect high-frequency areas. As a result, distortions in adversarial images are nearly invisible to the human eye, according to the contrast masking theory of human vision (Legge & Foley, 1980). For measuring the difference between attacked and original images/frames during a perturbation construction, we use adversarial mean absolute error in the frequency domain ($MAE^{*}$), which is calculated as follows:
$MAE^{*}=\frac{1}{HW}\displaystyle\sum_{i=0}^{(H-1)}\sum_{j=0}^{(W-1)}|I^{a*}_{ij}-I_{ij}^{*}|,$ (3)
@@ 3.2 One-iteration attack
Figure 1: An overview of the proposed IOI adversarial attack.
@@ 3.3 Frequency module
The frequency module extracts features from the original $I$ and perturbed $I^{p}$ images for further processing. It decomposes them into low-frequency (LF) and high-frequency (HF) components using the Fast Fourier Transform (FFT) and Inverse Fast Fourier Transform (IFFT). The LF component keeps fundamental content information, while the HF component contains noise and texture details. A threshold $tr^{*}$ for dividing frequencies in fhighest and (1-f)lowest coefficients is used to calculate indexes $d_{f}(I)$ of the highest FFT coefficient of the original image $I$ (Equation 6).
@@
We use $d_{f}(I)$ to extract LF and HF components from the original image, as shown in Equation 7. These components are further utilized in the weighting module to enhance the visual quality of adversarial video frames. HF components of the perturbed image $I^{p}$ are extracted using indexes $d_{f}(I)$ from the original image. Thus, the perturbations will be transferred into HF areas of the original image (Equation 8). The parameter $f$ allows fine-tuning the visibility of perturbations and relative attack gain at different levels. Lower values of $f$ result in more visible perturbations within the frequency module and higher relative gain.
@@ 3.4 Weighting module
Various image processing applications (Lin et al., 2005; Liu et al., 2010) operate under the assumption that distortions in low-variance regions are more visible to the human eye than in high-variance areas. The weighting module generates a weighting map for the input image based on the variance map. A similar approach was used in previous studies (Croce & Hein, 2019; Karli et al., 2021) for generating adversarial examples. We introduce additional features to enhance the weights obtained from the variance map method. The proposed weights map generation method is described in Equation 9. Initially, we compute local variance and local mean maps for an image $x$, determining standard deviations and means for both axes using a window size of 3 for each colour channel. Next, we derive the relative local variance of the image by dividing the local variance map by the local mean map and normalize the weights to the range [0, 1]. In the next step, we zero out 1% and compute the square root of the weights.
@@ 3.5 Mathematical properties
Proof. Each element of the $I^{a}-I$ are estimated using Equation 10 and representation of high-frequency FFT components as re-transformed 2D discrete Fourier transform without fof highest frequencies. The complete proof is presented in Appendix A.
The statement above demonstrates that the proposed method guarantees theoretical restriction of $l_{\infty}$ norm of the adversarial image, which depends on initial attack strength expressed in $MAE^{*}$ and parameter $f$ for truncating frequencies. Higher $f$ lead to saving more frequencies and providing better visual quality of the generated IOI adversarial image/video.
@@ 4 Experiments
Datasets. NIPS2017 image dataset (2017) was used to evaluate attacks on three NR metrics. It includes 1,000 images of a 299$\times$299 resolution. For evaluating methods on videos, we used 12 videos with 1280$\times$720 resolution from the DERF dataset (2001). Descriptions of these videos are available in the Appendix E. We extracted 75 frames from each original video and saved an attacked video with a frame rate of 25 frames per second, resulting in 3-second videos. The datasets licences allow usage for research purposes.
@@
Methodology. To evaluate the efficiency of adversarial attacks, we considered three factors: relative gain (Equation 4), speed and the objective or subjective quality of perturbed images/videos. To compare these three factors, we fixed two of them by aligning the speed and relative gain of all methods and compared the objective and subjective visual quality of adversarial images/videos. Since most of the attack time is spent on backpropagation, we executed each attack for one iteration to standardize the speed of all attacks. We also conducted additional experiments with multiple iterations, results presented in the Appendix C. In one-iteration comparison, we combined the results of Zhang et al. and SSAH attacks with FGSM attacks in all tables because one iteration of these attacks is equivalent to one iteration of FGSM. Zhang et al. use an FR quality metric to preserve image fidelity, and SSAH uses the distance between low-frequency information of two images. Since the distorted image used in both methods appears only during the first iteration, these fidelity-preserving components yield zero gradients.
@@
We conducted a crowd-sourced subjective comparison using Subjectify.us (sub, Accessed: Jan 2024) to get subjective scores for adversarial videos. Original and adversarial videos were compressed using the x264 video codec with a CRF value 16 (preset “Medium”). Each participant was asked to choose the video of the superior visual quality from a random pair of videos shown sequentially. An option “Can’t choose” was also available for them. Videos were pre-downloaded in the browser to prevent delays in playback, and participants had the flexibility to replay the videos multiple times. Each participant compared 12 video pairs; two of the 12 pairs were special verification pairs with apparent differences in visual quality. Answers from 200 participants who failed the verification were excluded. We collected 8220 responses from 685 participants who passed verification and calculated subjective scores using the Bradley-Terry model (Bradley & Terry, 1952). More details about the subjective experiment setup are presented in Appendix F.
@@ 5 Results
The proposed IOI method showed higher SSIM, VIF and LPIPS scores for all attacked NR metrics. The PSNR score of our attack method is lower than that of other methods, which means that IOI changes more information in images; however, the perturbations are hidden in the images’ texture/contrast regions. AdvJND method showed promising results for attacking Hyper-IQA and TReS models but failed on PaQ-2-PiQ. NVW performed well on PaQ-2-PiQ and TReS.
Table 3 contains the results of objective comparison on 12 videos from the DERF dataset (2001) and attacking the PaQ-2-PiQ (Ying et al., 2020) metric. We applied attacks with greater intensity on videos to enhance distinguishability for further subjective comparison, so the average relative gain was 15.3%. The objective results are similar to the comparison on the images: IOI outperformed prior methods on SSIM, VIF and LPIPS metrics and showed comparable PSNR scores. The per-video results are in the Appendix G.
Subjective comparison. The subjective scores obtained from pairwise comparisons showed that the IOI attack generates adversarial videos of better visual quality: it holds a quality of 2.97, while other methods’ scores are below 2.16. Confidence intervals for IOI and other methods do not intersect. The intervals intersect for some prior methods since their adversarial videos are noisy and flickering, making their subjective quality difficult to rank. The video with the highest distinguishability was “Blue Sky” with tree branches swaying against the smooth sky. The video with the lowest distinguishability was “Rush Hour” with a highly noisy background. As shown in Figure 3, FGSM, NVW, and Korhonen et al. methods generate imperceptible distortion on the stone region but fail to suppress the perturbation on the sky background. AdvJND, UAP and FACPA cause visible distortions on the whole image. To produce the same relative gain as IOI using one iteration, $lr$ for other methods was high, yielding visible perturbations. As shown in Appendix C, all methods (except FGSM, UAP and FACPA) produce almost equivalent results at 20 iterations. But at one iteration, there is a crucial difference. The videos used for the subjective comparison are available in Supplementary Materials and link is hidden for a blind review.

### v2

@@ Abstract
No-reference image- and video-quality metrics are widely used in video processing benchmarks. The robustness of learning-based metrics under video attacks has not been widely studied. In addition to having success, attacks on metrics that can be employed in video processing benchmarks must be fast and imperceptible. This paper introduces an Invisible One-Iteration (IOI) adversarial attack on no-reference image and video quality metrics. The proposed method uses two modules to ensure high visual quality and temporal stability of adversarial videos and runs for one iteration, which makes it fast. We compared our method alongside eight prior approaches using image and video datasets via objective and subjective tests. Our method exhibited superior visual quality across various attacked metric architectures while maintaining comparable attack success and speed. We made the code available on GitHub: https://github.com/katiashh/ioi-attack.
@@ 1 Introduction
NR metrics are employed in various image- and video-processing benchmarks, such as super-resolution (PIRM-SR, 2019; Khrulkov & Babenko, 2021; Ma et al., 2017), video generation (VideoGeneration, 2024), video compression (Ghadiyaram et al., 2017). For some tasks, NR metrics show even better performance than FR metrics; for example, video super-resolution (SR-MSU, 2023) or video compression by new encoding standards such as H.266/VVC (Antsiferova et al., 2022). The developers of video-processing algorithms can integrate adversarial attacks on quality metrics into their methods to achieve higher positions in public benchmarks. Today, such cheating can be detected in benchmarks that publish subjective comparisons along with objective ones. For example, in MSU Codec Comparison 2021 (CC-MSU, 2021), the leaderboard by a learning-based metric VMAF, which was shown to be vulnerable to attacks (Siniukov et al., 2022; Zvezdakova et al., 2019), differs from a subjective one. Algorithms in these benchmarks compete in both visual quality and speed, and a high speed is essential for some real-life applications like universal encoding for video compression (from one to ten frames per second). Thus, to cheat in video processing benchmarks, it is profitable for an attacker to inject an imperceptible perturbation into the video without significantly decreasing the method speed.
@@
    High speed of an attack. The attack must operate at high speed for practical viability. A slow attack holds limited practical significance, as its integration into video processing algorithms will greatly slow down the method and lower its position in benchmarks.
@@
Our research aims to investigate the potential of injecting fast, invisible and temporally consistent adversarial attacks on NR quality metrics. This paper introduces the Invisible One-Iteration (IOI) adversarial attack for images and videos. To achieve high attack speed, our method yields perturbation by calculating the gradient of an attacked model using one access to the model. We further show that a one-iteration attack for each frame is more efficient than a many-iteration attack applied to only some frames. To keep the temporal stability of a perturbed video, we make our attack invisible to the human eye by using weighting and frequency modules. The proposed attack operates in a white-box scenario, which does not limit its applicability for benchmarks: usually, a comparison methodology is known, and quality metrics are published for reproducibility. The primary contributions of this work can be summarized as follows:
@@
Our code is publicly available at https://github.com/katiashh/ioi-attack.
@@ 2 Related Work
Numerous studies showed that $L_{p}$ norms are not suitable as a distance metric to evaluate perceptual image quality (Sharif et al., 2018; Fezza et al., 2019; Wang et al., 2004; Johnson et al., 2016; Isola et al., 2017). Perturbations in images generated under $L_{p}$ norm constraints often result in noisy pixels within smooth areas of the original image, which is easily perceptible to the human eye. Several adversarial attacks prioritize the visual quality of the generated adversarial images and use more sophisticated restrictions on perturbations than the bare implementation of $L_{p}$ constraints.
Zhang et al. (2020) introduces a novel approach AdvJND by incorporating just noticeable difference (JND) (Yang et al., 2005) coefficients into the $L_{\infty}$ norm constraint during adversarial example generation. These coefficients account for the human eye’s ability to perceive the threshold of changes in an image. The authors employed I-FGSM and FGSM algorithms as baselines. To enhance the visual fidelity of adversarial images, they amplified original perturbations obtained from the FGSM or I-FGSM methods by scaled JND coefficients.
SSAH (Luo et al., 2022) adversarial attack targets two objectives: semantic similarity of images and low-frequency constraint. The first component is crafted with a focus on image classification tasks. The second component adds perturbations within high-frequency regions by minimizing the difference between low-frequency information of adversarial and original images. Usually, it requires many iterations to converge and produce good visual quality.
Korhonen et al. (2022b) introduced an iterative attack method targeting NR quality metrics, employing a Sobel filter to hide distortions within textured regions. In each iteration, the model’s gradient under attack concerning the input is multiplied by a spatial activity map derived from the original image via the Sobel filter. This map highlights areas with substantial texture, enhancing the visual fidelity of the resulting adversarial images.
Zhang et al. (2022a) introduced an iterative adversarial image crafting approach leveraging various FR metrics like Chebyshev distance, SSIM, LPIPS, and DISTS to manage visual distortions. Their method involved the iterative minimization of a loss function consisting of two components: the attacked loss and a loss based on some differential FR image quality metric.
Karli et al. (2021) introduced the Normalized Variance Weighting (NVW) method aimed at amplifying perturbations within high-variance regions of images. This technique can be utilized alongside gradient attacks like FGSM or I-FGSM. Additionally, the authors proposed the LPIPS-minimization method, aiming to enhance perceptual quality by minimizing the LPIPS distance between the original and adversarial images while ensuring the classifier remains deceived. However, this minimization method is exclusively effective for discrete tasks, such as classification or detection, while applying it to quality metrics will lead to eliminating relative gain.
@@ 3.1 Problem formulation
The adversarial attack on the NR quality metric $M$ is usually modelled as follows:
@@
To ensure a high visual quality of the perturbed image, instead of using $l_{p}$ norms that are inefficient for this task, we formulate the problem as follows:
$\begin{array}[]{cc}\underset{I^{a}}{\mathrm{argmax}}\{M(I^{a})-M(I)\},\\ \frac{L_{f}(I^{a})\cdot L_{f}(I)}{||L_{f}(I^{a})||_{2}||L_{f}(I)||_{2}}\geq 1-\epsilon^{*},\end{array}$ (2)
where $L_{f}$ is a low-frequency filter, $\epsilon^{*}$ is a small constant. When low-frequency components of the clear and adversarial images are closely aligned, the perturbations mainly affect high-frequency areas. As a result, distortions in adversarial images are nearly invisible to the human eye, according to the contrast masking theory of human vision (Legge & Foley, 1980). For measuring the difference between attacked and original images/frames during a perturbation construction, we use adversarial mean absolute error in the frequency domain ($MAE^{*}$), which is calculated as follows:
$MAE^{*}(I^{a},I)=\frac{1}{HW}\displaystyle\sum_{i=0}^{(H-1)}\sum_{j=0}^{(W-1)}|I^{a*}_{ij}-I_{ij}^{*}|,$ (3)
@@ 3.2 One-iteration attack
Figure 1: An overview of the proposed IOI adversarial attack. $I$ is stands for input image, $I^{p}$ – FGSM attacked image and $I^{a}$ – the final IOI attacked image. Weights map is calculated using formula 9.
@@ 3.3 Frequency module
The frequency module extracts features from the original $I$ and perturbed $I^{p}$ images for further processing. It decomposes them into low-frequency (LF) and high-frequency (HF) components using the Fast Fourier Transform (FFT) and Inverse Fast Fourier Transform (IFFT). The LF component keeps fundamental content information, while the HF component contains noise and texture details. A threshold $tr^{*}$ for dividing frequencies in fhighest and (1-f)the lowest coefficients is used to calculate indexes $d_{f}(I)$ of the highest FFT coefficient of the original image $I$ (Equation 6).
@@
We use $d_{f}(I)$ to extract LF and HF components from the original image, as shown in Equation LABEL:eq:freq32. These components are further utilized in the weighting module to enhance the visual quality of adversarial video frames. HF components of the perturbed image $I^{p}$ are extracted using indexes $d_{f}(I)$ from the original image. Thus, the perturbations will be transferred into HF areas of the original image (Equation LABEL:eq:freq31). The parameter $f$ allows fine-tuning the visibility of perturbations and relative attack gain at different levels. Lower values of $f$ result in more visible perturbations within the frequency module and higher relative gain.
@@ 3.4 Weighting module
Various image processing applications (Lin et al., 2005; Liu et al., 2010) operate under the assumption that distortions in low-variance regions are more visible to the human eye than in high-variance areas. The weighting module generates a weighting map for the input image based on the variance map. A similar approach was used in previous studies (Croce & Hein, 2019; Karli et al., 2021) for generating adversarial examples. We introduce additional features to enhance the weights obtained from the variance map method. The proposed weights map generation method is described in Equation 9. Initially, we compute local variance and local mean maps for an image $x$, determining standard deviations and means for both axes using a window size of 3 for each colour channel. Next, we derive the relative local variance of the image by dividing the local variance map by the local mean map and normalizing the weights to the range [0, 1]. In the next step, we zero out 1% and compute the square root of the weights.
@@ 3.5 Mathematical properties
Proof. Each element of the $(I^{a}-I)$ is estimated using Equation 10 and representation of high-frequency FFT components as re-transformed 2D discrete Fourier transform without fof highest frequencies. The complete proof is presented in the Appendix A.
The statement above demonstrates that the proposed method guarantees theoretical restriction of $l_{\infty}$ norm of the adversarial image, which depends on initial attack strength expressed in $MAE^{*}$ and parameter $f$ for truncating frequencies. Higher $f$ leads to saving more frequencies and providing better visual quality of the generated IOI adversarial image/video.
@@ 4 Experiments
Datasets. NIPS2017 image dataset (2017) was used to evaluate attacks on three NR metrics. It includes 1,000 images of a 299$\times$299 resolution. For evaluating methods on videos, we used 12 videos with 1280$\times$720 resolution from the DERF dataset (2001). Descriptions of these videos are available in the Appendix J. We extracted 75 frames from each original video and saved an attacked video with a frame rate of 25 frames per second, resulting in 3-second videos. The datasets licenses allow usage for research purposes.
@@
Methodology. To evaluate the efficiency of adversarial attacks, we considered three factors: relative gain (Equation 4), speed and the objective or subjective quality of perturbed images/videos. To compare these three factors, we fixed two of them by aligning the speed and relative gain of all methods and compared the objective and subjective visual quality of adversarial images/videos. Since most of the attack time is spent on backpropagation, we executed each attack for one iteration to standardize the speed of all attacks. We also conducted additional experiments with multiple iterations, results presented in the Appendix C. In a one-iteration comparison, we combined the results of Zhang et al. and SSAH attacks with FGSM attacks in all tables because one iteration of these attacks is equivalent to one iteration of FGSM. Zhang et al. use an FR quality metric to preserve image fidelity, and SSAH uses the distance between low-frequency information of two images. Since the distorted image used in both methods appears only during the first iteration, these fidelity-preserving components yield zero gradients.
@@
We conducted a crowd-sourced subjective comparison using Subjectify.us (sub, Accessed: Jan 2024) to get subjective scores for adversarial videos. Original and adversarial videos were compressed using the x264 video codec with a CRF value 16 (preset “Medium”). Each participant was asked to choose the video of the superior visual quality from a random pair of videos shown sequentially. An option “Can’t choose” was also available for them. Videos were pre-downloaded in the browser to prevent delays in playback, and participants had the flexibility to replay the videos multiple times. Each participant compared 12 video pairs; two of the 12 pairs were special verification pairs with apparent differences in visual quality. Answers from 200 participants who failed the verification were excluded. We collected 8220 responses from 685 participants who passed verification and calculated subjective scores using the Bradley-Terry model (Bradley & Terry, 1952). More details about the subjective experiment setup are presented in the Appendix K.
@@ 5 Results
The proposed IOI method showed higher SSIM, VIF, and LPIPS scores for all attacked NR metrics. The PSNR score of our attack method is lower than that of other methods, which means that IOI changes more information in images; however, the perturbations are hidden in the images’ texture/contrast regions. AdvJND method showed promising results for attacking Hyper-IQA and TReS models but failed on PaQ-2-PiQ. NVW performed well on PaQ-2-PiQ and TReS.
Table 3 contains the results of objective comparison on 12 videos from the DERF dataset (2001) and attacking the PaQ-2-PiQ (Ying et al., 2020) metric. We applied attacks with greater intensity on videos to enhance distinguishability for further subjective comparison, so the average relative gain was 15.3%. The objective results are similar to the comparison on the images: IOI outperformed prior methods on SSIM, VIF, and LPIPS metrics and showed comparable PSNR scores. The per-video results are in the Appendix L.
Subjective comparison. The subjective scores obtained from pairwise comparisons showed that the IOI attack generates adversarial videos of better visual quality: it holds a quality of 2.97, while other methods’ scores are below 2.16. Confidence intervals for IOI and other methods do not intersect. The intervals intersect for some prior methods, since their adversarial videos are noisy and flickering, making their subjective quality difficult to rank. The video with the highest distinguishability was “Blue Sky” with tree branches swaying against the smooth sky. The video with the lowest distinguishability was “Rush Hour” with a highly noisy background. As shown in Figure 3, FGSM, NVW, and Korhonen et al. methods generate imperceptible distortion on the stone region but fail to suppress the perturbation on the sky background. AdvJND, UAP and FACPA cause visible distortions on the whole image. To produce the same relative gain as IOI using one iteration, $lr$ for other methods, was high, yielding visible perturbations. As shown in the Appendix C, all methods (except FGSM, UAP and FACPA) produce almost equivalent results at 20 iterations. But at one iteration, there is a crucial difference. The videos used for the comparison are available at https://github.com/katiashh/ioi-attack.